OpenStudy (anonymous):

In ∫ 2x(x^(2)+3)dx, if we let u=x^(2)+3,then 1.the integral form is u^(n)du and du=2x 2.the integral form u^(n)du cannot be used because n=1 3.the integral form is u^(n)du and du=2dx 4.the integral form is u^(n)du and du=2xdx

OpenStudy (anonymous):

\[here\int\limits2x(x ^{2}+3)dx_{}^{}\] for simplification we will take (x^2+3)=u by diff. both side wrt x we get 2x+0=du/dx so du = 2xdx & integral form is udu. i think u r missing sum thing in the question that it must be integration of 2x(x^(2)+3)^n.dx is not it........